OpenStudy (anonymous):

what is the value of a1 for a geometric sequence with a4=40 and a6=160 a-2 b-5 c-8 d-10

jimthompson5910 (jim_thompson5910):

Hints: a5 = a4*r a6 = a5*r ---> a6 = (a4*r)*r

jimthompson5910 (jim_thompson5910):

a6 = (a4*r)*r turns into a6 = a4*r^2

jimthompson5910 (jim_thompson5910):

because a6 = a4*r^2, a4=40 and a6=160, we can say a6 = a4*r^2 160 = 40*r^2 r = ???

OpenStudy (anonymous):

What does r stand for

jimthompson5910 (jim_thompson5910):

r is the common ratio

OpenStudy (anonymous):

duh

OpenStudy (anonymous):

so the common ratio is what?

jimthompson5910 (jim_thompson5910):

if you have the fourth term a4, you multiply that fourth term by r to get the fifth term this applies in general to any term (not just the 4th term)

jimthompson5910 (jim_thompson5910):

here's an example of a geometric sequence 1, 2, 4, 8, 16, 32 Notice how each term is multiplied by 2 to get the next one (eg: 4 ---> 8, 8 = 4*2)

jimthompson5910 (jim_thompson5910):

so in my example, r = 2

OpenStudy (anonymous):

Ok so the common ratio in my problem is 2?

jimthompson5910 (jim_thompson5910):

no, that's for my example

jimthompson5910 (jim_thompson5910):

I don't know what your common ratio r value is yet

OpenStudy (anonymous):

I know, i realized i made a mistake when i posted that one

jimthompson5910 (jim_thompson5910):

you have to solve 160 = 40*r^2

OpenStudy (anonymous):

How do i find the common ratio for that

OpenStudy (anonymous):

nevermind

jimthompson5910 (jim_thompson5910):

Solving 160 = 40*r^2 gives you r = ???

OpenStudy (anonymous):

wouldnt it be 2?

jimthompson5910 (jim_thompson5910):

yeah it's actually 2 it's a coincidence that I used 2 as well

OpenStudy (anonymous):

Ok thats what i thought haha

jimthompson5910 (jim_thompson5910):

Now we use the formula \[\Large a_{n} = a*(r)^{n-1}\]

jimthompson5910 (jim_thompson5910):

plug in n = 4, r = 2 to get \[\Large a_{n} = a*(r)^{n-1}\] \[\Large a_{4} = a*(2)^{4-1}\] Then plug in \(\large a_{4} = 40\) to get \[\Large 40 = a*(2)^{4-1}\] now solve for 'a'

OpenStudy (anonymous):

Ok now i can plug in one of a b c or d

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

you could, but solving is more direct

jimthompson5910 (jim_thompson5910):

I recommend doing it this way because you shouldn't be dependent on multiple choice. What if this was a free response or fill in the blank problem?

OpenStudy (anonymous):

Good point

jimthompson5910 (jim_thompson5910):

I guess though, having multiple choice there means you should take advantage of it to save time. So you should know how to do it both ways.

OpenStudy (anonymous):

So its b

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

that's 'a' or a1

OpenStudy (anonymous):

Sweet thanks

jimthompson5910 (jim_thompson5910):

you're welcome